Overnight, Kyle tried cycling our fridges. One of them cycled successfully, but the other didn't -- and after looking at the data, it appears it started too warm for a successful fridge cycle given that we tried it a day earlier than we have in the past after having filled helium. So we're going to try again tonight, with hopefully better results. Still, the results were good enough that Kevin and Kate could get some of the preliminary bolo info needed to get the system tuned.
In the morning, we also had a visit from McMurdo's helicopter crew to go over various options for recovering the cryostat in the event that the experiment lands close enough to be reached by helicopter. Shaul also finished cleaning the primary mirror, a task that took him the better part of a full working day. In the meantime, I was already scheduled to cut foam for our inner frame baffling, so I was tasked with making a cover for the mirror. So I busted out my trusty high school geometry skills (and refreshed them with wikipedia's article on ellipses) and got two screws and some string and drew an awesome ellipse to match the outline of the primary mirror (which has an elliptical rim but looks like a circle from the point of view of the incoming light to the telescope).
In the afternoon, KyleH, Jeff, and I cut a LOT of foam. Matt and Franky set about debugging the few SQUIDS (5 out of 112) that had some problems tuning by opening up the cryostat RF can and poking about. Seth, Britt, and Michele spent much of the afternoon debugging some spontaneous flight computer rebooting issues. They tracked down the issue to some flaky power pins inside the computer crate and are implementing a solution.
In the other high bay, BLAST-Pol took their gondola outside so they could do some beam maps with their telescope and a chopped source on a shipping container a few hundred feet away from the high bay.